A parallelogram is formed in R3 by the vectors = (3, 2, –3) and = (4, 1, 5). The point P = (0, 2, 3). a. Determine the location of the vertices. b. Determine the vectors representing the diagonals. c. Determine the length of the diagonals.
1
Expert's answer
2019-11-25T12:52:50-0500

Solution:a) Let PMNK - a parallelogram.


"\overrightarrow {a}=\overrightarrow {PK}"

"\overrightarrow {PK} (x; y-2; z-3)"


"x=3; y-2=2; y=4; z-3=-3; z=0. K (3;4;0)"



"\overrightarrow {b}= \overrightarrow {PM}"


"\overrightarrow {PM} (x; y-2; z-3);"




"x=4; y-2=1; y=3; z-3=5; z=8. M(4;3;8)"


"\overrightarrow {a} = \overrightarrow {MN}"



"\overrightarrow {MN} (x-4; y-3; z-8);"


"x-4=3; x=7; y-3=2; y=5; z-8=-3; z=5. N(7;5;5)"


b)



"\overrightarrow {PN} (7-0;5-2;5-3); \overrightarrow {PN} (7;3;2)"


"\overrightarrow {MK} (3-4;4-3;0-8); \overrightarrow {MK} (-1;1;-8)"



c)


"\vert \overrightarrow {PN} \vert= \sqrt {49+9+4} = \sqrt {62}"

"\vert \overrightarrow {MK} \vert = \sqrt {1+1+64} = \sqrt {66}"